OpenStudy (anonymous):

The measure of angle 1 is 5 less than 4 times the measure of angle 2. if angle 1 and angle 2 form a linear pair, what are there measures?

OpenStudy (anonymous):

Sum of angles in linear pair is 180 Angle 2 = x Angle 1 = 4x -5 x+4x -5 = 180 5x -5 = 180 5x = 185 Angle 2 = x =185/5 = 37 Angle 1 = 4x -5 = 4*37 -5 = 148 -5 = 143

OpenStudy (dumbcow):

Let A be angle1, B be angle 2 A = 4B-5 A+B = 180 By substitution (4B-5)+B = 180 5B = 185 B = 37 A = 180-37 = 143

OpenStudy (anonymous):

thank you guys so much!